# Dynamic Diffusion Schrödinger Bridge in Astrophysical Observational Inversions

### v2

@@ 4.2 Prior methods and evaluation setup
The conventional methods involve astrostatistical techniques to infer physical distributions with simplified physical assumptions. For molecular density prediction, existing popular techniques include statistical conversions such as 2-component (2PLF) and 3-component power-law fitting (3PLF). For magnetic field, Davis–Chandrasekhar–Fermi (DCF) method (Davis, 1951; Chandrasekhar and Fermi, 1953; Beck, 2016) has been the classic method for estimating the magnetic field strength, which relies on the assumption of an equipartition between the magnetic field energy and the turbulent kinetic energy of the gas.
Machine learning methods.
@@
Recent studies have also adopted both generative and discriminative machine learning approaches. Discriminative models are typically convolutional neural networks (CNNs) trained with reconstruction losses (Xu et al., 2020a, b). Generative Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) were recently introduced as an alterantive.
@@
We compare with both conventional astrostatistical and ML-based methods. Specifically for ML, in addition to the state-of-the-art conditional DDPMs, we also re-implement a U-Net (Ronneberger et al., 2015) from CASI-2D (Xu et al., 2020a) as a pixel-to-pixel approach for comparison. Notably, while recent works suggest that generative models offer improved sample quality in image synthesis, their advantage in scientific prediction, especially under physically grounded distribution shift (OOD), remains unclear. Our study explicitly tests this hypothesis by comparing probabilistic and pixel-to-pixel modeling under controlled simulation setups with similar model scales.
Evaluation metrics.
@@
Our evaluation adheres to established protocols in astrophysical studies, emphasizing the distributional patterns of relative error between predictions and ground truth values. Specifically, we assess the normalized relative error (e.g., $\Delta\rho/\rho$), where an ideal prediction should produce a distribution centered around zero. Although this distribution may not strictly follow a Gaussian form, a symmetric shape with a narrow spread around zero indicates accurate and well-calibrated predictions. In contrast, distributions with large bias or heavy tails reflect systematic under- or over-estimation. To quantitatively capture these characteristics, we report the mean $\mu$ (gt - pred) and standard deviation $\sigma$ of the relative error distribution as primary metrics. These values provide insights into the central tendency and variability of model performance across different test cases.
@@
Implementation details.
To ensure a fair comparison among the tested ML methods, we fix the U-Net backbone architecture across the discriminative U-Net, conditional DDPMs, and our proposed Astro-DSB, resulting in approximately 81M learnable parameters for each model. For the U-Net baseline, we train the model for 20 epochs using Mean Square Error (MSE) loss. For conditional DDPMs, we follow the setup from prior works (Xu et al., 2023a, 2025) and adopt a 1000-step schedule with the standard variational lower bound loss (Ho et al., 2020). Our proposed Astro-DSB is trained with the same time discretization (1000 steps) using the objective defined in Eq. 5. All models are trained with a batch size of 16 using the AdamW optimizer with a learning rate of 5e-5, running on an AWS cluster with 4 NVIDIA T4 GPUs. Notably, while conditional DDPMs require around 400 epochs to converge, Astro-DSB converges in only around 100 epochs under the same setup, reducing overall training time by approximately 75%. This aligns with our expectations and observations in prior vision DSB works, as the target distribution $p_{1}$ now lies closer to the physically meaningful observation distribution $p_{0}$, compared to the case where $p_{0}$ is a zero-mean Gaussian in DDPMs. During inference, the runtime per testing case ranges from approximately 6 to 30 seconds, depending on whether the step skipping (Song et al., 2021) is applied. Notably, we observe no significant difference in prediction performance across these sampler settings in our work.
@@
Methods $\mu$ $p_{\Delta\rho/\rho}$ ($\downarrow$|) $\sigma$ $p_{\Delta\rho/\rho}$ ($\downarrow$|) ME $\mu$ $p_{\Delta\rho/\rho}$ ($\downarrow$|) $\sigma$ $p_{\Delta\rho/\rho}$ ($\downarrow$|) ME
2PLF -2.50$\pm$7E-4 7.11$\pm$2E-3 -715.95$\pm$1.88 -5.77$\pm$3E-4 7.21$\pm$1E-3 -1680.82$\pm$0.74
3PLF -0.77$\pm$5E-4 4.44$\pm$2E-3 -715.78$\pm$1.88 -3.23$\pm$2E-4 4.21$\pm$7E-4 -1680.60$\pm$0.61
U-NET -0.25$\pm$1E-4 1.29$\pm$4E-4 118.24$\pm$0.94 5.01$\pm$6E-4 7.08$\pm$1E-3 1313.60$\pm$0.61
cDDPMs -0.05$\pm$6E-5 0.61$\pm$6E-4 4.61$\pm$1.05 2.88$\pm$5E-4 5.33$\pm$7E-4 1050.51$\pm$0.61
Astro-DSB (Ours) -0.02$\pm$1E-4 0.71$\pm$3E-4 32.09$\pm$0.65 0.51$\pm$2E-4 2.32$\pm$3E-3 -5.04$\pm$1.32
@@ 4.3 Experimental results and analysis
We present the quantitative evaluation results for density and magnetic field prediction in Tab. 1 and Tab. 2, respectively. Our proposed Astro-DSB consistently achieves the best or second-best results across both ID and OOD settings, particularly demonstrating robust superiority in OOD predictions, as indicated by the lower mean ($\mu$) and standard deviation ($\sigma$) of the relative error distribution. In contrast, traditional astrostatistical methods such as power-law fitting (PLF) and DCF exhibit large error values, while U-Net and DDPMs suffer from more severe performance degradation under OOD conditions.
In addition, the training cost of Astro-DSB is only 25% of the previous SOTA conditional DDPMs method.
@@
Figure 3: Multiple qualitative visualizations for relative error comparisons. Results from Astro-DSB present more balanced and unbiased prediction error patterns compared to other ML methods, particularly for OOD testing cases. Best viewed in color and with zoom-in.
@@
Methods $\mu$ $p_{\Delta\mathcal{B}/\mathcal{B}}$ ($\downarrow$|) $\sigma$ $p_{\Delta\mathcal{B}/\mathcal{B}}$ ($\downarrow$|) ME $\mu$ $p_{\Delta\mathcal{B}/\mathcal{B}}$ ($\downarrow$|) $\sigma$ $p_{\Delta\mathcal{B}/\mathcal{B}}$ ($\downarrow$|) ME
DCF 11.80$\pm$8E-4 21.78$\pm$4E-4 507.74$\pm$0.03 5.19$\pm$8E-4 9.64$\pm$3E-3 137.76$\pm$0.03
PLF -0.11$\pm$5E-5 1.23$\pm$1E-4 -1.04$\pm$1E-3 1.03$\pm$9E-5 1.47$\pm$1E-4 17.14$\pm$2E-3
U-NET -0.06$\pm$2E-5 0.38$\pm$5E-5 154.76$\pm$0.08 -0.32$\pm$6E-5 0.74$\pm$7E-5 146.38$\pm$0.02
cDDPMs 0.07$\pm$2E-5 0.38$\pm$6E-5 123.99$\pm$0.04 0.12$\pm$7E-5 0.72$\pm$2E-4 150.00$\pm$0.04
Astro-DSB (Ours) 0.15$\pm$2E-5 0.55$\pm$5E-5 235.85$\pm$0.08 0.19$\pm$6E-5 0.67$\pm$9E-5 131.89$\pm$0.02
@@
Method $\mu$ $p_{\Delta\rho/\rho}$ ($\downarrow$|) $\sigma$ $p_{\Delta\rho/\rho}$ ($\downarrow$|) ME $\mu$ $p_{\Delta\rho/\rho}$ ($\downarrow$|) $\sigma$ $p_{\Delta\rho/\rho}$ ($\downarrow$|) ME
w/o M1 & M2 -0.19$\pm$3E-4 1.75$\pm$0.04 66.54$\pm$1.49 4.24$\pm$2E-3 14.71$\pm$0.12 1086.66$\pm$0.66
w/o M1 -3E-3$\pm$7E-5 0.65$\pm$6E-4 35.86$\pm$0.44 -1.84$\pm$5E-4 5.06$\pm$4E-3 -2297.48$\pm$1.10
w/o M2 7E-3$\pm$1E-4 1.63$\pm$5E-4 37.15$\pm$1.53 -14.66$\pm$5E-3 54.95$\pm$3.47 -12069.44$\pm$9.44
Astro-DSB (Full) -0.02$\pm$1E-4 0.71$\pm$3E-4 32.09$\pm$0.65 0.51$\pm$2E-4 2.32$\pm$3E-3 -5.04$\pm$1.32
Extended robustness on observations from Taurus B213.
We further test our Astro-DSB on the Herschel column density map of Taurus B213 (Li and Goldsmith, 2012; Palmeirim et al., 2013) using the model trained purely on synthetic simulations. Fig. 4 visualizes the observed column density map alongside our predicted volume density map. Despite the lack of pixel-level ground truth in real-world observations, our predictions demonstrate a coherent spatial structure and effectively capture the filamentary features of B213, as qualitatively verified by astrophysicists. Notably, the results align well with single-pointing constraints derived from indirect measurements using the density tracer HC3(Li and Goldsmith, 2012), which suggests that Astro-DSB can generalize effectively beyond simulated scenarios, providing reliable physical predictions even on real observational data. More experimental results are in the appendices.
@@
Ablation studies on key methodological components.
We evaluate the effectiveness of two key components in our Astro-DSB design: M1: Noise alignment and M2: Observable enhancement, by systematically removing them and reporting the results in Tab. 3. Our findings reveal that the absence of either module significantly degrades performance, with M2 being particularly effective for cases of OOD and stable prediction indicated as indicated by $\sigma$ values. The full Astro-DSB achieves the best balance of prediction accuracy and generalization robustness.
@@ 5 Conclusion and Discussion
In this work, we address a class of astrostatistical inverse prediction tasks by introducingAstro-DSB, a dynamic generative modeling framework based on the Diffusion Schrödinger Bridge. Targeting key physical quantities within Giant Molecular Clouds, Astro-DSB learns the probabilistic transition between partial observables and true physical states, offering a more flexible alternative to conventional one-step or purely discriminative approaches. Our study reveals that, in contrast to static tasks where dynamic modeling may appear excessive, explicitly learning distributional transitions is both meaningful and beneficial in physically grounded systems demonstrated via improved generalization in out-of-domain simulations and real-world observables. Overall, Astro-DSB offers better alignment with physical interpretability and supports the broader potential of probabilistic modeling in scientific discovery.
Limitation and future direction.
Despite the improved robustness and performance to unseen simulation conditions, our method does not explicitly model the full forward physical equations that govern the GMC evolution. While this is reasonable given the intractable nature of observational inversion problems in this work, future extensions could involve integrating multiple intermediate states, leveraging richer supervision signals, or directly encoding the dynamic physical laws into the generative modeling process to further strengthen the alignment between ML and physical dynamics for extended applications with well-defined physical forward processes.

### v3

@@ 4.2 Comparing methods and evaluations
Astrostatistical methods. The conventional methods involve astrostatistical techniques to infer physical distributions with simplified physical assumptions. For molecular density prediction, existing popular techniques include statistical conversions such as 2-component (2PLF) and 3-component power-law fitting (3PLF). For magnetic field, Davis–Chandrasekhar–Fermi (DCF) method (Davis, 1951; Chandrasekhar and Fermi, 1953; Beck, 2016) has been the classic method for estimating the magnetic field strength, which relies on the assumption of an equipartition between the magnetic field energy and the turbulent kinetic energy of the gas.
@@
Machine learning methods. Recent studies have also adopted machine learning techniques, which can be broadly categorized into discriminative and generative approaches. Discriminative models such as convolutional neural networks (CNNs) trained with reconstruction losses have been widely adopted (Xu et al., 2020a, b). From the generative modeling perspective, conditional Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) have emerged as a recent popular alternative to tackle those observational prediction problems.
@@
In our experiments, we compare with both conventional and ML-based methods. Specifically for ML, in addition to the state-of-the-art conditional DDPMs, we also re-implement a U-Net (Ronneberger et al., 2015) from CASI-2D (Xu et al., 2020a) as a pixel-to-pixel approach for comparison. Notably, while recent works suggest that generative models offer improved sample quality in image synthesis, their advantage in scientific prediction, especially under physically grounded distribution shift (OOD), remains unclear. Our study explicitly tests this hypothesis by comparing probabilistic and pixel-to-pixel modeling under controlled simulation setups with similar model scales.
@@
Evaluations. Our evaluation adheres to established protocols in astrophysical studies, emphasizing the distributional patterns of relative error between predictions and ground truth values. Specifically, we assess the normalized relative error (e.g., $\Delta\rho/\rho$), where an ideal prediction should produce a distribution centered around zero. Although this distribution may not strictly follow a Gaussian form, a symmetric shape with a narrow spread around zero indicates accurate and well-calibrated predictions. In contrast, distributions with large bias or heavy tails reflect systematic under- or over-estimation. To quantitatively capture these characteristics, we report the mean $\mu$ (gt - pred) and standard deviation $\sigma$ of the relative error distribution as primary metrics. These values provide insights into the central tendency and variability of model performance across different test cases.
@@
Implementation details. To ensure a fair comparison among the tested ML methods, we fix the U-Net backbone architecture across the discriminative U-Net, conditional DDPMs, and our proposed Astro-DSB, resulting in approximately 81M learnable parameters for each model. For the U-Net baseline, we train the model for 20 epochs using Mean Square Error (MSE) loss. For conditional DDPMs, we follow the setup from prior works (Xu et al., 2023a, 2025) and adopt a 1000-step schedule with the standard variational lower bound loss (Ho et al., 2020). Our proposed Astro-DSB is trained with the same time discretization (1000 steps) using the objective defined in Eq. 5. All models are trained with a batch size of 16 using the AdamW optimizer with a learning rate of 5e-5, running on an AWS cluster with 4 NVIDIA T4 GPUs. Notably, while conditional DDPMs require around 400 epochs to converge, Astro-DSB converges in only around 100 epochs under the same setup, reducing overall training time by approximately 75%. This aligns with our expectations and observations in prior vision DSB works, as the target distribution $p_{1}$ now lies closer to the physically meaningful observation distribution $p_{0}$, compared to the case where $p_{0}$ is a zero-mean Gaussian in DDPMs. During inference, the runtime per testing case ranges from approximately 6 to 30 seconds, depending on whether the step skipping (Song et al., 2021) is applied. Notably, we observe no significant difference in prediction performance across these sampler settings in our work.
@@
Methods $\mu$ - $p_{\Delta\rho/\rho}$ ($\downarrow|$) $\sigma$ - $p_{\Delta\rho/\rho}$ ($\downarrow|$) ME $\mu$ - $p_{\Delta\rho/\rho}$ ($\downarrow|$) $\sigma$ - $p_{\Delta\rho/\rho}$ ($\downarrow|$) ME
[…]
@@ 4.3 Experimental results and analysis
We present the quantitative evaluation results for density and magnetic field prediction in Tab. 1 and Tab. 2, respectively. Our proposed Astro-DSB consistently achieves the best or second-best results across both ID and OOD settings, particularly demonstrating robust superiority in OOD predictions, as indicated by the lower mean ($\mu$) and standard deviation ($\sigma$) of the relative error distribution. In contrast, traditional astrostatistical methods such as power-law fitting (PLF) and DCF exhibit large error values, while U-Net and DDPMs suffer from more severe performance degradation under OOD conditions. In particular, the training cost of Astro-DSB is only 25% of the previous SOTA conditional DDPMs method.
@@
Figure 3: Multiple qualitative visualizations for relative error comparisons. (a) 2D histogram showing the distribution of ground truth (x-axis) versus predictions (y-axis). An accurate and unbiased prediction lies along the 1-to-1 diagonal line. (b) Relative error between the ground truth and predictions. (c) Residual maps showing the difference between model predictions and the ground truth across different methods. Results from Astro-DSB present more balanced and unbiased prediction error patterns compared to other ML methods, particularly for OOD testing cases. Best viewed in color and with zoom-in.
@@
Methods $\mu$ - $p_{\Delta\mathcal{B}/\mathcal{B}}$ ($\downarrow|$) $\sigma$ - $p_{\Delta\mathcal{B}/\mathcal{B}}$ ($\downarrow|$) ME $\mu$ - $p_{\Delta\mathcal{B}/\mathcal{B}}$ ($\downarrow|$) $\sigma$ - $p_{\Delta\mathcal{B}/\mathcal{B}}$ ($\downarrow|$) ME
[…]
@@
Method $\mu$ - $p_{\Delta\rho/\rho}$ ($\downarrow|$) $\sigma$ - $p_{\Delta\rho/\rho}$ ($\downarrow|$) ME $\mu$ - $p_{\Delta\rho/\rho}$ ($\downarrow|$) $\sigma$ - $p_{\Delta\rho/\rho}$ ($\downarrow|$) ME
[…]
Extended robustness on observations from Taurus B213. We further test our Astro-DSB on the Herschel column density map of Taurus B213 (Li and Goldsmith, 2012; Palmeirim et al., 2013) using the model trained purely on synthetic simulations. Fig. 4 visualizes the observed column density map alongside our predicted volume density map. Despite the lack of pixel-level ground truth in real-world observations, our predictions demonstrate a coherent spatial structure and effectively capture the filamentary features of B213, as qualitatively verified by astrophysicists. Notably, the results align well with single-pointing constraints derived from indirect measurements using the density tracer HC3(Li and Goldsmith, 2012), which suggests that Astro-DSB can generalize effectively beyond simulated scenarios, providing reliable physical predictions even on real observational data. More experimental results are in the appendices.
@@
Ablation studies on key methodological components. We evaluate the effectiveness of two key components in our Astro-DSB design: M1: Noise alignment and M2: Observable enhancement, by systematically removing them and reporting the results in Tab. 3. Our findings reveal that the absence of either module significantly degrades performance, with M2 being particularly effective for cases of OOD and stable prediction indicated as indicated by $\sigma$ values. The full Astro-DSB achieves the best balance of prediction accuracy and generalization robustness.
@@ 5 Conclusion and Discussion
In this work, we address a class of astrostatistical inverse prediction tasks by introducingAstro-DSB, a dynamic generative modeling framework based on the Diffusion Schrödinger Bridge. Targeting key physical quantities within GMCs, Astro-DSB learns the probabilistic transition between partial observables and true physical states, offering a more flexible alternative to conventional one-step or purely discriminative approaches. Our study reveals that, in contrast to static tasks where dynamic modeling may appear excessive, explicitly learning distributional transitions is both meaningful and beneficial in physically grounded systems demonstrated via improved generalization in out-of-domain simulations and real-world observables. Overall, Astro-DSB offers better alignment with physical interpretability and supports the broader potential of probabilistic modeling in scientific discovery.
Limitation and future direction. Despite the improved robustness and performance to unseen simulation conditions, our method does not explicitly model the full forward physical equations that govern the GMC evolution. While this is reasonable given the intractable nature of observational inversion problems in this work, future extensions could involve integrating multiple intermediate states, leveraging richer supervision signals, or directly encoding the dynamic physical laws into the generative modeling process to further strengthen the alignment between ML and physical dynamics for extended applications with well-defined physical forward processes.